# Controlling Continuous Relaxation for Combinatorial Optimization

### v3

@@ Abstract
Unsupervised learning (UL)-based solvers for combinatorial optimization (CO) train a neural network whose output provides a soft solution by directly optimizing the CO objective using a continuous relaxation strategy. These solvers offer several advantages over traditional methods and other learning-based methods, particularly for large-scale CO problems. However, UL-based solvers face two practical issues: (I) an optimization issue where UL-based solvers are easily trapped at local optima, and (II) a rounding issue where UL-based solvers require artificial post-learning rounding from the continuous space back to the original discrete space, undermining the robustness of the results. This study proposes a Continuous Relaxation Annealing (CRA) strategy, an effective rounding-free learning method for UL-based solvers. CRA introduces a penalty term that dynamically shifts from prioritizing continuous solutions, effectively smoothing the non-convexity of the objective function, to enforcing discreteness, eliminating the artificial rounding. Experimental results demonstrate that CRA significantly enhances the performance of UL-based solvers, outperforming existing UL-based solvers and greedy algorithms in complex CO problems. It also effectively eliminates the artificial rounding and accelerates the learning.
@@ 1 Introduction
The objective of combinatorial optimization (CO) problems is to find the optimal solution from a discrete space, and these problems are fundamental in many real-world applications (Papadimitriou and Steiglitz, 1998). Most CO problems are NP-hard or NP-complete, making it challenging to solve large-scale problems within feasible computational time. Traditional methods frequently depend on heuristics to find approximate solutions; however, considerable insights into the specific problems are required. Recently, several studies have used machine learning methods to handle CO problems by learning heuristics. Most of these studies focus on supervised learning (SL)-based solvers (Hudson et al., 2021; Joshi et al., 2019; Gasse et al., 2019; Selsam et al., 2018; Khalil et al., 2016), which require optimal solutions to the CO problem as supervision during training. However, optimal solutions are hard to get in practice, and SL-based solvers often do not generalize well (Yehuda et al., 2020). Reinforcement learning (RL)-based solvers (Yao et al., 2019; Chen and Tian, 2019; Yolcu and Póczos, 2019; Nazari et al., 2018; Khalil et al., 2017; Bello et al., 2016) do not require such optimal solutions but frequently suffer from notoriously unstable training due to poor gradient estimation and hard explorations (Mnih et al., 2015; Tang et al., 2017; Espeholt et al., 2018). Recently, unsupervised learning (UL)-based solvers have attracted much attention (Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). UL-based solvers follow a continuous relaxation approach. They train a UL model to output a “soft” solution to the relaxed CO problem by directly optimizing the differentiable objective function, offering significantly stable and fast training even for large-scale CO problems. Notably, physics-inspired GNN (PI-GNN) solver (Schuetz et al., 2022a) employs graph neural networks (GNN) to automatically learn specific heuristics for a single instance and performs on par with or outperforms existing solvers for CO problems with millions of variables without optimal solutions.
While offering some advantages over traditional and other machine learning-based solvers, they face two practical issues. The first issue is an optimization issues where UL-based solvers are easily trapped at local optima. Angelini and Ricci-Tersenghi (2023) demonstrated that the PI-GNN solver could not achieve results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) in maximum independent set (MIS) problems on random regular graphs (RRG). Wang and Li (2023) also pointed out the importance of using dataset/history and initialization of GNN using output of greedy solvers for PI-GNN solver to overcome the optimization issues. This issue is a crucial bottleneck for its applicability in various real-world applications. The second issue relates to the inherent ambiguity in the continuous relaxation approach. This approach requires artificial rounding from the soft solution, which may include continuous values, back to the original discrete solution, potentially undermining the robustness of the results. Indeed, while linear relaxation can provide an optimal solution for original discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically leads to solutions with $\nicefrac{{1}}{{2}}$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974), where existing rounding methods (Schuetz et al., 2022b; Wang et al., 2022) completely lose their robustness.
To address these issues, we propose the Continuous Relaxation Annealing (CRA) strategy. The CRA introduces a penalty term to control the continuity and discreteness of the relaxed variables, with a parameter $\gamma$ to regulate the intensity of this penalty term. If the parameter $\gamma$ is small, the relaxed variable tends to favor continuous solutions. In contrast, a large $\gamma$ value leads the relaxed variables to favor the discrete space. This penalty term also effectively eliminates local optimum. Moreover, a small $\gamma$ forces the loss function to approach a simple convex function, encouraging active exploration in the continuous space. The CRA also includes an annealing process, where $\gamma$ is increased progressively until the relaxed variables are nearly discrete values, eliminating the artificial rounding from the continuous to the original discrete space after learning. In this study, the solver that applies the CRA to the PI-GNN solver is referred to as the CRA-PI-GNN solver.
@@
We also demonstrate the benefits of the CRA through experiments on benchmark CO problems, including MIS, maximum cut (MaxCut), and diverse bipartite matching (DBM) problems across graphs of varying sizes and degrees. The experimental results show that the CRA significantly enhances the performance of the PI-GNN solver, outperforming the original PI-GNN solver, other state-of-the-art learning-based baselines, and greedy algorithms. This improvement is achieved by directly optimizing each instance without any history/dataset. Additionally, these experiments indicate that the CRA accelerates the learning process of the PI-GNN solver. Notably, these results overcome the limitations pointed out by Angelini and Ricci-Tersenghi (2023); Wang and Li (2023), highlighting the further potential of UL-based solvers.
@@
We use the shorthand expression $[N]=\{1,2,\ldots,N\}$,  $N\in{\mathbb{N}}$. Here, $I_{N}\in{\mathbb{R}}^{N\times N}$ denotes an $N\times N$ identity matrix, ${\bm{1}}_{N}$ denotes the vector $(1,\ldots,1)^{\top}\in{\mathbb{R}}^{N}$, and ${\bm{0}}_{N}$ denotes the vector $(0,\ldots,0)^{\top}\in{\mathbb{R}}^{N}$. $G(V,E)$ denotes an undirected graph, where $V$ is the set of nodes with cardinality $|V|=N$, and $E\subseteq V\times V$ is the set of edges. For a graph $G(V,E)$, $A_{ij}$ denote the adjacency matrix with $A_{ij}=0$ if an edge $(i,j)$ does not exist and $A_{ij}>0$ if the edge connects.
@@ 2 Background
where $C\in{\mathcal{C}}$ denotes instance parameters, e.g., a graph $G=(V,E)$, where $V$ is the node set, and $E$ is the edge set, and $f:{\mathcal{X}}\times{\mathcal{C}}\to{\mathbb{R}}$ represents the cost function. In addition, ${\bm{x}}=(x_{i})_{1\leq i\leq N}\in\{0,1\}^{N}$ is a binary vector to be optimized, and ${\mathcal{X}}(C)\subseteq\{0,1\}^{N}$ is the space of feasible solutions, typically specified by the following equality and inequality constraints,
${\mathcal{X}}(C)=\{{\bm{x}}\in\{0,1\}^{N}\mid\forall i\in[I],~{}g_{i}({\bm{x}};C)\leq 0,~{}\forall j\in[J],~{}h_{j}({\bm{x}};C)=0\}.$
Here, for $i\in[I]$, $g_{i}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the inequality constraint, and for $j\in[J]$, $h_{j}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the equality constraint. Following UL-based solvers (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020), we first rewrite the constrained problem into an equivalent unconstrained form via the penalty method (Smith et al., 1997):
@@
Here, for all $i\in[I+J]$, $v_{i}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the penalty term, which increases the value when the constraints are violated, e.g., the penalty term is defined as follows:
$\forall i\in[I],j\in[J],~{}v_{i}({\bm{x}};C)=\max(0,g_{i}({\bm{x}};C)),~{}~{}\forall j\in[J],~{}v_{j}({\bm{x}};C)=(h_{j}({\bm{x}};C))^{2},$
and ${\bm{\lambda}}=(\lambda_{i})_{1\leq i\leq I+J}\in{\mathbb{R}}^{I+J}$ represents the penalty parameters that control the tradeoff between satisfying the constraints and optimizing the cost. Note that the penalty for violations of the constraints becomes more significant as ${\bm{\lambda}}$ increases. In the following, we show the example of this formulation.
@@
The MIS problem is a fundamental NP-hard problem (Karp, 2010) defined as follows. Given an undirected graph $G(V,E)$, an independent set (IS) is a subset of nodes ${\mathcal{I}}\in V$ where any two nodes in the set are not adjacent. The MIS problem attempts to find the largest IS, which is denoted ${\mathcal{I}}^{\ast}$. In this study, $\rho$ denotes the IS density, where $\rho=|{\mathcal{I}}|/|V|$. To formulate the problem, a binary variable $x_{i}$ is assigned to each node $i\in V$. Then the MIS problem is formulated as follows:
@@
where the first term attempts to maximize the number of nodes assigned $1$, and the second term penalizes the adjacent nodes marked $1$ according to the penalty parameter $\lambda$.
@@ 2.1 Unsupervised learning based solvers
Learning for CO problems involve leaning an algorithm ${\mathcal{A}}_{{\bm{\theta}}}(\cdot):{\mathcal{C}}\to\{0,1\}^{N}$ parameterized by a nural network (NN), where ${\bm{\theta}}$ denotes the parameters. Given an instance parameter $C\in{\mathcal{C}}$, this algorithm generates a valid solution $\hat{{\bm{x}}}={\mathcal{A}}_{\theta}(C)\in{\mathcal{X}}(C)$ and minimizes $f(\hat{{\bm{x}}};C)$. Several approaches have been proposed to learn ${\mathcal{A}}_{\theta}$. This study focuses on UL-based solvers, where no labeled solution ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}\in{\mathcal{X}}(C)}f({\bm{x}};C)$ is employed during training (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). In the following, we describe the details of the UL-based solvers. The UL-based solvers employ continuous relaxation strategy for directly training neural networks (NN). The continuous relaxation strategy involves reformulating an original CO problem into a continuous optimization problem by converting discrete variables into continuous variables. A typical example of continuous relaxation can be expressed as follows:
@@
where ${\bm{p}}=(p_{i})_{1\leq i\leq N}\in[0,1]^{N}$ represents a set of relaxed continuous variables, i.e., each binary variable $x_{i}\in\{0,1\}$ is relaxed to a continuous one $p_{i}\in[0,1]$, and $\hat{f}:[0,1]^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the relaxation of $f$, satisfying $\hat{f}({\bm{x}};C)=f({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. The relation between the constraint $v_{i}$ and its relaxation $\hat{v}_{i}$ is similar for $i\in[I+J]$, i.e., $\forall i\in[I+J],~{}\hat{v}_{i}({\bm{x}};C)=v_{i}({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. Wang et al. (2022) and Schuetz et al. (2022a) formulated ${\mathcal{A}}_{\theta}(C)$ as the relaxed continuous variables, i.e., ${\mathcal{A}}_{\theta}(\cdot):{\mathcal{C}}\to[0,1]^{n}$. In the following discussions, to represent the parametrization of the relaxed variables explicitly, we denote ${\mathcal{A}}_{\theta}$ by ${\bm{p}}_{\theta}$. Then, ${\bm{p}}_{\theta}$ is optimized by minimizing the following label-independent function directly:
@@
After training the relaxed solution ${\bm{p}}_{{\bm{\theta}}}$ is transformed into discrete variables by artificial rounding ${\bm{p}}_{{\bm{\theta}}}$, such that $\forall i\in[N],~{}x_{i}=\mathrm{int}(p_{{\bm{\theta}},i}(C))$ using a threshold (Schuetz et al., 2022a), or via a greedy method (Wang et al., 2022). Based on this formulation, two types of schemes for UL-based solvers have been developed:
@@
One approach, proposed by Karalias and Loukas (2020), aims to automatically learn typically effective heuristics from history/dataset instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$ and then apply these learned heuristics to new instance $C^{\ast}$ through inference. Note that this method assumes that the training dataset can be easily obtained or that meaningful data augmentation is possible. Specifically, given a set of training instances ${\mathcal{D}}=(C_{\mu})$, IID sampled from a distribution $P(C)$, the goal is to minimize the averaged loss function $\min_{{\bm{\theta}}}\sum_{\mu=1}^{P}l({\bm{\theta}};C_{\mu},{\bm{\lambda}})$. However, this method cannot guarantee quality for a test instance $C^{\ast}$. Even if the training instances ${\mathcal{D}}$ are numerous and the test instance $C$ also follows $P(C)$, a low averaged performance ${\mathbb{E}}_{C\sim P(C)}[\hat{l}(\theta;C)]$ may not guarantee a low $l(\theta;C)$ for on specific $C$. To address this issues, Wang and Li (2023) introduced a meta-learning approach, where NNs aim for good initialization for future instances rather than direct solutions.
@@
Another approach, referred to as the PI-GNN solver (Schuetz et al., 2022a, b), automatically learns instance-specific heuristics for a given single instance using the instance parameter $C$ by directly employing Eq. (4). This approach studied CO problems on graphs, i.e., $C=G(V,E)$, and employed GNNs for the relaxed variables $p_{\theta}(G)$. An $L$-layered GNN is trained to directly minimize $\hat{l}({\bm{\theta}};C,{\bm{\lambda}})$ in Eq. (4), taking as input a graph $G$ and the embedding vectors on the nodes, and outputting the relaxed solution ${\bm{p}}_{{\bm{\theta}}}(G)\in[0,1]^{N}$. A detailed description of GNNs is given in Appendix E.2. Note that this setting is applicable even when the training dataset ${\mathcal{D}}$ is not easily obtainable. However, learning to minimize Eq. 4 for a single instance can be time-consuming compared with the inference process. Nonetheless, for large-scale CO problems, Schuetz et al. (2022a, b) demonstrated superiority over other solvers in terms of both time and solution performance. This study focuses on the advancement of this UL based solver for a single instance.

### v4

@@ Abstract
Unsupervised learning (UL)-based solvers for combinatorial optimization (CO) train a neural network that generates a soft solution by directly optimizing the CO objective using a continuous relaxation strategy. These solvers offer several advantages over traditional methods and other learning-based methods, particularly for large-scale CO problems. However, UL-based solvers face two practical issues: (I) an optimization issue, where UL-based solvers are easily trapped at local optima, and (II) a rounding issue, where UL-based solvers require artificial post-learning rounding from the continuous space back to the original discrete space, undermining the robustness of the results. This study proposes a Continuous Relaxation Annealing (CRA) strategy, an effective rounding-free learning method for UL-based solvers. CRA introduces a penalty term that dynamically shifts from prioritizing continuous solutions, effectively smoothing the non-convexity of the objective function, to enforcing discreteness, eliminating artificial rounding. Experimental results demonstrate that CRA significantly enhances the performance of UL-based solvers, outperforming existing UL-based solvers and greedy algorithms in complex CO problems. Additionally, CRA effectively eliminates artificial rounding and accelerates the learning process.
@@ 1 Introduction
The objective of combinatorial optimization (CO) problems is to find the optimal solution from a discrete space, and these problems are fundamental in many real-world applications (Papadimitriou and Steiglitz, 1998). Most CO problems are NP-hard or NP-complete; making it challenging to solve large-scale problems within feasible computational time. Traditional methods frequently depend on heuristics to find approximate solutions; however, considerable insights into the specific problems are required. Alternatively, problems can be formulated as integer linear programming (ILP) and solved using ILP solvers. However, ILP lacks scalability when applied to problems with graph structures.
@@
Recently, several studies have used machine learning methods to handle CO problems by learning heuristics. Most of these studies focus on supervised learning (SL)-based solvers (Hudson et al., 2021; Joshi et al., 2019; Gasse et al., 2019; Selsam et al., 2018; Khalil et al., 2016), which require optimal solutions to CO problems as supervision during training. However, obtaining optimal solutions is challenging in practice, and SL-based solvers often fail to generalize well (Yehuda et al., 2020). Reinforcement learning (RL)-based solvers (Yao et al., 2019; Chen and Tian, 2019; Yolcu and Póczos, 2019; Nazari et al., 2018; Khalil et al., 2017; Bello et al., 2016) avoid the need for optimal solutions but often suffer from notoriously unstable training due to poor gradient estimation and hard explorations (Mnih et al., 2015; Tang et al., 2017; Espeholt et al., 2018). Unsupervised learning (UL)-based solvers have recently attracted much attention (Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). UL-based solvers follow a continuous relaxation approach, training a UL model to output a soft solution to the relaxed CO problem by directly optimizing a differentiable objective function, offering significantly stable and fast training even for large-scale CO problems. Notably, the physics-inspired GNN (PI-GNN) solver (Schuetz et al., 2022a) employs graph neural networks (GNN) to automatically learn instance-specific heuristics and performs on par with or outperforms existing solvers for CO problems with millions of variables without optimal solutions. While these offer some advantages over traditional and other machine learning-based solvers, they face two practical issues. The first issue is an optimization issues where UL-based solvers are easily trapped at local optima. Due to this issue, Angelini and Ricci-Tersenghi (2023) demonstrated that the PI-GNN solver (Schuetz et al., 2022a) could not achieve results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) on maximum independent set (MIS) problems in random regular graphs (RRG). Wang and Li (2023) also pointed out the importance of using dataset or history, and initializing the GNN with outputs from greedy solvers to help the PI-GNN solver overcome optimization challenges. This issue is a crucial bottleneck to the applicability of this method across various real-world applications. The second issue relates to the inherent ambiguity of the continuous relaxation approach. This approach necessitates artificial rounding from the soft solution, which may include continuous values, back to the original discrete solution, potentially undermining the robustness of the results. While linear relaxation can provide an optimal solution for original discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically leads to solutions with $1/2$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974), in which existing rounding methods (Schuetz et al., 2022b; Wang et al., 2022) completely lose their robustness. For NP-hard problems with graph structures, such as the MIS and MaxCut, semidefinite programming (SDP) relaxations have been proposed as effective approximation methods (Lovász, 1979; Goemans and Williamson, 1995). However, these approaches rely on rounding techniques, such as spectral clustering (Von Luxburg, 2007), to transform relaxed solutions into feasible ones, which often fails to obtain optimal solutions.
@@
To address these issues, we propose the Continuous Relaxation Annealing (CRA). CRA introduces a penalty term to control the continuity and discreteness of the relaxed variables, with a parameter $\gamma$ to regulate the intensity of this penalty term. When the parameter $\gamma$ is small, the relaxed variable tends to favor continuous solutions, whereas a large $\gamma$ biases them toward discrete values. This penalty term also effectively eliminates local optimum. Moreover, a small $\gamma$ forces the loss function to approach a simple convex function, encouraging active exploration within the continuous space. CRA also includes an annealing process, where $\gamma$ is gradually increased until the relaxed variables approach discrete values, eliminating the artificial rounding from the continuous to the original discrete space after learning. In this study, the solver that applies the CRA to the PI-GNN solver is referred to as the CRA-PI-GNN solver. We also demonstrate the benefits of the CRA through experiments on benchmark CO problems, including MIS, maximum cut (MaxCut), and diverse bipartite matching (DBM) problems across graphs of varying sizes and degrees. The experimental results show that the CRA significantly enhances the performance of the PI-GNN solver, outperforming the original PI-GNN solver, other state-of-the-art learning-based baselines, and greedy algorithms. This improvement is achieved by directly optimizing each instance without any history, e.g., previous optimal solutions and the information of other instances. Additionally, these experiments indicate that the CRA accelerates the learning process of the PI-GNN solver. Notably, these results overcome the limitations pointed out by Angelini and Ricci-Tersenghi (2023); Wang and Li (2023), highlighting the further potential of UL-based solvers.
@@
We use the shorthand expression $[N]=\{1,2,\ldots,N\}$, where $N\in{\mathbb{N}}$. $I_{N}\in{\mathbb{R}}^{N\times N}$ denotes an $N\times N$ identity matrix, ${\bm{1}}_{N}$ denotes the vector $(1,\ldots,1)^{\top}\in{\mathbb{R}}^{N}$, and ${\bm{0}}_{N}$ denotes the vector $(0,\ldots,0)^{\top}\in{\mathbb{R}}^{N}$. $G(V,E)$ represents an undirected graph, where $V$ is the set of nodes with cardinality $|V|=N$, and $E\subseteq V\times V$ denotes the set of edges. For a graph $G(V,E)$, $A_{ij}$ denotes the adjacency matrix, where $A_{ij}=0$ if an edge $(i,j)$ does not exist and $A_{ij}>0$ if the edge is present.
@@ 2 Background
where $C\in{\mathcal{C}}$ denotes instance-specific parameters, such as a graph $G=(V,E)$, and ${\mathcal{C}}$ represents the set of all possible instances. $f:{\mathcal{X}}\times{\mathcal{C}}\to{\mathbb{R}}$ denotes the cost function. Additionally, ${\bm{x}}=(x_{i})_{1\leq i\leq N}\in\{0,1\}^{N}$ is a binary vector to be optimized, and ${\mathcal{X}}(C)\subseteq\{0,1\}^{N}$ denotes the feasible solution space, typically defined by the following equality and inequality constraints.
${\mathcal{X}}(C)=\{{\bm{x}}\in\{0,1\}^{N}\mid\forall i\in[I],~{}g_{i}({\bm{x}};C)\leq 0,~{}\forall j\in[J],~{}h_{j}({\bm{x}};C)=0\},~{}~{}I,J\in{\mathbb{N}},$
Here, for $i\in[I]$, $g_{i}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ denotes the inequality constraint, and for $j\in[J]$, $h_{j}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ denotes the equality constraint. Following UL-based solvers (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020), we reformulate the constrained problem into an equivalent unconstrained form using the penalty method (Smith et al., 1997):
@@
Here, for all $i\in[I+J]$, $v:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the penalty term, which increases when the constraints are violated. For example, the penalty term is defined as follows:
$\forall i\in[I],j\in[J],~{}v_{i}({\bm{x}};C)=\max(0,g_{i}({\bm{x}};C)),~{}~{}\forall j\in[J],~{}v_{j}({\bm{x}};C)=(h_{j}({\bm{x}};C))^{2}$
and ${\bm{\lambda}}=(\lambda_{i})_{1\leq i\leq I+J}\in{\mathbb{R}}^{I+J}$ denotes the penalty parameters that control the trade-off between constraint satisfaction and cost optimization. Note that, as ${\bm{\lambda}}$ increases, the penalty for constraint violations becomes more significant. In the following, we provide an example of this formulation.
@@
The MIS problem is a fundamental NP-hard problem (Karp, 2010), defined as follows. Given an undirected graph $G(V,E)$, an independent set (IS) is a subset of nodes ${\mathcal{I}}\in V$ where any two nodes are not adjacent. The MIS problem aims to find the largest IS, denoted as ${\mathcal{I}}^{\ast}$. In this study, $\rho$ denotes the IS density, defined as $\rho=|{\mathcal{I}}|/|V|$. Following Schuetz et al. (2022a), a binary variable $x_{i}$ is assigned to each node $i\in V$. The MIS problem can be formulated as follows:
@@
where the first term maximizes the number of nodes assigned a value of $1$, and the second term penalizes adjacent nodes assigned $1$ according to the penalty parameter $\lambda$.
@@ 2.1 Unsupervised learning based solvers
Learning for CO problems involves training an algorithm ${\mathcal{A}}_{{\bm{\theta}}}(\cdot):{\mathcal{C}}\to\{0,1\}^{N}$ parameterized by a neural network (NN), where ${\bm{\theta}}$ denotes the parameters. For a given instance $C\in{\mathcal{C}}$, this algorithm generates a valid solution $\hat{{\bm{x}}}={\mathcal{A}}_{\theta}(C)\in{\mathcal{X}}(C)$ and aims to minimize $f(\hat{{\bm{x}}};C)$. Several approaches have been proposed to train ${\mathcal{A}}_{\theta}$. This study focuses on UL-based solvers, which do not use a labeled solution ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}\in{\mathcal{X}}(C)}f({\bm{x}};C)$ during training (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). In the following, we outline the details of the UL-based solvers.
The UL-based solvers employ a continuous relaxation strategy to train NN. This continuous relaxation strategy reformulates a CO problem into a continuous optimization problem by converting discrete variables into continuous ones. A typical example of continuous relaxation is expressed as follows:
@@
where ${\bm{p}}=(p_{i})_{1\leq i\leq N}\in[0,1]^{N}$ represents a set of relaxed continuous variables, where each binary variable $x_{i}\in\{0,1\}$ is relaxed to a continuous counterpart $p_{i}\in[0,1]$, and $\hat{f}:[0,1]^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ denotes the relaxed form of $f$ such that $\hat{f}({\bm{x}};C)=f({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. The relation between each constraint $v_{i}$ and its relaxation $\hat{v}_{i}$ is similar for $i\in[I+J]$, meaning that $\forall i\in[I+J],~{}\hat{v}_{i}({\bm{x}};C)=v_{i}({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. Wang et al. (2022) and Schuetz et al. (2022a) formulated ${\mathcal{A}}_{\theta}(C)$ as the relaxed continuous variables, defined as ${\mathcal{A}}_{\theta}(\cdot):{\mathcal{C}}\to[0,1]^{n}$. In the following discussions, we denote ${\mathcal{A}}_{\theta}$ as ${\bm{p}}_{\theta}$ to make the parametrization of the relaxed variables explicit. Then, ${\bm{p}}_{\theta}$ is optimized by directly minimizing the following label-independent function:
@@
After training, the relaxed solution ${\bm{p}}_{{\bm{\theta}}}$ is converted into discrete variables using artificial rounding ${\bm{p}}_{{\bm{\theta}}}$, where $\forall i\in[N],~{}x_{i}=\mathrm{int}(p_{{\bm{\theta}},i}(C))$ based on a threshold (Schuetz et al., 2022a), or alternatively, a greedy method (Wang et al., 2022). Two types of schemes for UL-based solvers have been developed based on this formulation.
@@
One approach, proposed by Karalias and Loukas (2020), aims to automatically learn effective heuristics from historical dataset instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$ and then apply these learned heuristics to a new instance $C^{\ast}$, through inference. Note that this method assumes that either the training dataset is easily obtainable or that meaningful data augmentation is feasible. Specifically, given a set of training instances ${\mathcal{D}}=(C_{\mu})$, sampled independently and identically from a distribution $P(C)$, the goal is to minimize the average loss function $\min_{{\bm{\theta}}}\sum_{\mu=1}^{P}l({\bm{\theta}};C_{\mu},{\bm{\lambda}})$. However, this method does not guarantee quality for a test instance, $C^{\ast}$. Even if the training instances ${\mathcal{D}}$ are extensive and the test instance $C$ follows $P(C)$, low average performance ${\mathbb{E}}_{C\sim P(C)}[\hat{l}(\theta;C)]$ may not guarantee a low $l(\theta;C)$ for on a specific $C$. To address this issue, Wang and Li (2023) introduced a meta-learning approach where NNs aim to provide good initialization for future instances rather than direct solutions.
@@
Another approach, known as the PI-GNN solver (Schuetz et al., 2022a, b), automatically learns instance-specific heuristics for a single instance using the instance parameter $C$ by directly applying Eq. (4). This approach addresses CO problems on graphs, where $C=G(V,E)$, and employs GNNs for the relaxed variables $p_{\theta}(G)$. Here, an $L$-layered GNN is trained to directly minimize $\hat{l}({\bm{\theta}};C,{\bm{\lambda}})$, taking as input a graph $G$ and the embedding vectors on its nodes, and outputting the relaxed solution ${\bm{p}}_{{\bm{\theta}}}(G)\in[0,1]^{N}$. A detailed description of GNNs is provided in Appendix E.2. Note that this setting is applicable even when the training dataset ${\mathcal{D}}$ is difficult to obtain. The overparameterization of relaxed variables is expected to smooth the objective function by introducing additional parameters to the optimization problem, similar to the kernel method. However, minimizing Eq. 4 for a single instance can be time-consuming compared to the inference process. Nonetheless, for large-scale CO problems, this approach has been reported to outperform other solvers in terms of both computational time and solution quality (Schuetz et al., 2022a, b).
Note that, while both UL-based solvers for multiple instances (Type I) and individual instances (Type II) are valuable, this study focuses on advancing the latter: a UL-based solver for a single instance. Both types of solvers are applicable to cost functions that meet a particular requirement due to their reliance on a gradient-based algorithm to minimize Eq (4).
Assumption 2.1 (Differentiable cost function).
The relaxed loss function $\hat{l}({\bm{\theta}};C,{\bm{\lambda}})$ and its partial derivative $\nicefrac{{\partial\hat{l}({\bm{\theta}};C,{\bm{\lambda}})}}{{\partial{\bm{\theta}}}}$ are accessible during the optimization process.
These requirements encompass a nonlinear cost function and interactions involving many-body interactions, extending beyond simple two-body interactions.